OpenStudy (anonymous):

Evaluate the function at the given values of the independent variable and simplify. f(r^6)=15(r^6)/17|r^6|

OpenStudy (anonymous):

\[f(r) =15\frac{r^6}{17\left| r^6 \right|}\] find f(r^6) right?

OpenStudy (anonymous):

confirm the question because the way you wrote the question above is wrong

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

oops just noticed the error

OpenStudy (anonymous):

ok so all you have to do is plug in r^6, let me show you

OpenStudy (anonymous):

it supposed to say f(r^6)=15(x)/17|x|

OpenStudy (anonymous):

\[f(r^6) =15\frac{(r^6)^6}{17\left| (r^6)^6 \right|}\]

OpenStudy (anonymous):

ok wait ...

OpenStudy (anonymous):

so i know you plug in r^6 where there's an x. I just can't seem to get answer right on webwork

OpenStudy (anonymous):

ok now you changed the question completly

OpenStudy (anonymous):

so you are saying now this is the function f(x)=15(x)/17|x| find f(r^6) right! confirm please

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[f(r^6) =\frac{15r^6}{17\left| r^6 \right|}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so what is your final answer then?

OpenStudy (anonymous):

there is still one step left can you see it?

OpenStudy (anonymous):

no i got 15r6/17r6

OpenStudy (anonymous):

so cancle r6 then you would be left with?

OpenStudy (anonymous):

what is 10/10?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

what is r/r?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

and what is r6/r6?

OpenStudy (anonymous):

so its just 15/17?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hehehe well done

OpenStudy (anonymous):

lol, thanks you very much

OpenStudy (anonymous):

you're welcome so much